A jar contains 4 red marbles, 12 blue marbles and 6 green marbles. Three marbles are selected at random, with each marble being set aside after selection. What is the probability that at least 1 red marble is drawn?
a. 0.470
b. 0.338
c. 0.632
d. 0.530
e. 0.701

*DO NOT USE EXCEL*
Answer #1

A jar contains

4 red marbles

12 blue marbles

6 green marbles

Total 22 marbles

There are three marbles are selected from the jar

Probabality that atleast one red marble is drawn

= P( one red marble & two others) + P( two red marbles and one other) + P( all three red marbles )

= [4C1 * 18C2 / 22C3] + [ 4C2 * 18C1 / 22C3 ] + [ 4C3 *18C0 / 22C3 ]

=( 612 / 1540) + ( 108 / 1540) + ( 4 / 1540)

= 724 / 1540

= 0.470

Answer A) 0.470